Please s explain the below expression cos¢=9/15, Tan¢=?.

Grade:7

1 Answers

Hari Sundar Pyda
6 years ago
cos C=9/15We know that sin2C+ cos2C= 1.Sin2C= 1-cos2CSin2C= 1-(9/15)^2Sin2C= 1- 81/225Sin2C= 225-81/225Sin2C= 144/225SinC= 12/15TanC= sinC/cosC= 12/15 / 9/15TanC= 12/15*15/9TanC= 4/3.